OpenStudy (gavin39):

sSolve x − 2y = 6 3x − 6y = 18 (2, −2) (3, negative three halves) No Solutions Infinitely Many Solutions

OpenStudy (studygurl14):

x - 2y = 6 3x - 6y = 18 x = 6 + 2y Substitue it into the other equation. 3(6 + 2y) - 6y = 18 What do you do now to solve for y?

OpenStudy (gavin39):

Have to find the factors of each equation?

OpenStudy (studygurl14):

Well, solve for y... Do this first: 3(6 + 2y) - 6y = 18 18 + 6y - 6y = 18 What now?

OpenStudy (gavin39):

The equaton wuld have to be simplified 18 + 6y - 6y = 18 18 + 6y - 6y = 18

OpenStudy (studygurl14):

yes, but what next?

OpenStudy (gavin39):

Hmm not to sure to be honest

OpenStudy (studygurl14):

add/subtract like terms so.... what are like terms?